OpenStudy (anonymous):

Find the inverse function of f(x)=\sqrt(3 x+9). f^(-1)(x)=

OpenStudy (isaiah.feynman):

Procedure for finding inverse of a function. write the equation with respect to y switch x and y solve for y the new equation is the inverse.

OpenStudy (jmark):

Given f(x)=√(3x+9) then f^(-1)(x)=? let y=f(x) y=√(3x+9) y²=3x+9 y²-9/3=x there for x=(y²-9)/3 is the inverse function for f(x) Learn more about Inverse Secant Function at http://goo.gl/xYXYnM